# SLAM3R: Real-Time Dense Scene Reconstruction from Monocular RGB Videos

### v2

@@ Abstract
In this paper, we introduce SLAM3R, a novel and effective monocular RGB SLAM system for real-time and high-quality dense 3D reconstruction. SLAM3R provides an end-to-end solution by seamlessly integrating local 3D reconstruction and global coordinate registration through feed-forward neural networks. Given an input video, the system first converts it into overlapping clips using a sliding window mechanism. Unlike traditional pose optimization-based methods, SLAM3R directly regresses 3D pointmaps from RGB images in each window and progressively aligns and deforms these local pointmaps to create a globally consistent scene reconstruction - all without explicitly solving any camera parameters. Experiments across datasets consistently show that SLAM3R achieves state-of-the-art reconstruction accuracy and completeness while maintaining real-time performance at 20+ FPS. Code and weights at: https://github.com/PKU-VCL-3DV/SLAM3R.
@@ 1 Introduction
Figure 1: We introduce a novel dense SLAM system - SLAM3R. This system takes a monocular RGB video as input and reconstructs the scene as a dense pointcloud. The video is converted into short clips for local reconstruction (denoted as inner-window), which are then incrementally registered together (inter-window) to create a global scene model. This process runs in real-time, producing a reconstruction that is both accurate and complete.
@@
While monocular dense SLAM systems encounter the limitations mentioned earlier, recent advances in two-view geometry have shown promising potential. DUSt3R [62] introduces a purely end-to-end approach for learning dense reconstruction. Trained on large-scale datasets, its network is capable of producing high-quality dense reconstructions from paired images in real-time. However, for multiple views, a global optimization step is required to align these image pairs, which significantly hampers its efficiency. A concurrent work, Spann3R [59], extends DUSt3R to multi-view (video) scenarios through a pairwise incremental reconstruction pipeline. While this method accelerates the reconstruction process, it unfortunately results in significant accumulated drift and reduced reconstruction quality.
@@
  • We present a novel real-time system for end-to-end dense 3D reconstruction that directly predicts 3D pointmaps in a unified coordinate system.
@@ 2 Related Work
Traditional offline approaches.
Dense 3D pointcloud reconstruction is a long-standing problem in computer vision. Classical approaches to this problem first determine camera parameters using Structure from Motion (SfM) [51, 64, 45, 29, 31], followed by dense 3D points triangulation with Multi-View Stereo (MVS) [16, 46, 66, 58]. In recent years, neural implicit [34, 60, 28, 63, 8] and 3D Gaussian [17, 18, 11] representations have been applied to further enhance the quality of dense reconstruction. While these methods deliver high-quality results, they have a significant limitation: the requirement for offline processing to generate the final 3D model, which restricts their applicability in real-time scenarios. In this paper, we focus on online dense reconstruction in the context of Simultaneous Localization and Mapping (SLAM).
Dense SLAM.
Early works on SLAM [13, 4, 23, 14, 36, 35, 7] focused on reconstructing the structure of unknown environments while simultaneously localizing camera poses. These approaches prioritize real-time performance but produce only sparse structures of the scene. Dense SLAM approaches [37, 74, 5, 10, 25, 54, 15, 53, 20, 76, 9, 40] incorporate detailed scene geometry information to improve pose estimation. DROID-SLAM [54] introduces recurrent iterative updates of camera poses and pixel-wise depth estimates, while TANDEM [25] proposes an online MVS module for depth prediction. These systems enable real-time dense scene reconstruction. However, their focus on camera trajectory accuracy often results in incomplete and noisy 3D reconstruction. Neural implicit and Gaussian representations have also been integrated with dense SLAM systems [53, 20, 76, 40, 70, 30, 21, 65, 32, 19, 9, 40, 43, 42]. However, these approaches often rely on additional depth sensors or focus primarily on novel view synthesis rather than producing detailed geometric reconstruction.
@@
End-to-end dense 3D reconstruction.
DUSt3R [62] introduces the first purely end-to-end dense 3D reconstruction pipeline without reply on camera parameters. Recently, several works have adopted a similar approach for single-view reconstruction [61], feature matching [26], novel view synthesis [50, 68], and dynamics reconstruction [72]. These successes demonstrate the effectiveness of end-to-end dense point prediction, inspiring us to develop a dense reconstruction system with a similar methodology.
While DUSt3R operates in real-time for two-view predictions, its extension to multiple views involves exhaustive pairing images and performing an additional global optimization step. This process significantly increases computational time, thereby hindering its real-time performance. MASt3R [26] enhances the matching capability of DUSt3R by adding a match head, achieving more accurate keypoint correspondences for 3D reconstruction, but at the cost of increased computational time. More recently, the concurrent work Spann3R [59] extends DUSt3R with spatial memory. It takes a video as input and performs incremental scene reconstruction in a unified coordinate system without requiring global optimization. While this approach significantly improves runtime efficiency, the pairwise incremental reconstruction pipeline leads to noticeable accumulated drift. Unlike Spann3R, our networks at each hierarchy take multiple frames as input to minimize drift. Additionally, we propose a self-contained retrieval module that, when registering a new frame, this module selects not only its previous few frames but also other similar frames from long-term history for more global scene reference.
@@ 3 Method
Problem statement.
Given a monocular video consisting of a sequence of RGB image frames $\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$ that captures a static scene, the goal is to reconstruct its dense 3D poincloud $P\in\mathbb{R}^{M\times 3}$, where $M$ is the number of 3D points. Research in this field focuses on three key objectives: maximizing 3D points recovery for reconstruction completeness, improving the accuracy of each recovered point, and achieving these goals while preserving real-time performance.
System overview.
Figure 2 illustrates an overview of the proposed dense SLAM system. It consists of two main components: an Image-to-Points (I2P) network that recovers local 3D points from video clips, and a Local-to-World (L2W) network that registers local reconstructions into a global scene coordinate system. During the reconstruction of the dense point cloud, the system does not explicitly solve any camera parameters. Instead, it directly predicts 3D point maps in unified coordinate systems.
The system starts by applying a sliding window mechanism of length $L$ to convert the input video into short clips $\{W_{i}\in\mathbb{R}^{L\times H\times W\times 3}\}$. The I2P network then processes each window $W_{i}$ to recover local 3D pointmaps. Within each window, the system selects a keyframe to define a reference coordinate system for point reconstruction, as detailed in Sec. 3.1. By default, the stride of the sliding window is set to 1, ensuring each input frame in the video is selected at least once as a keyframe. For global scene reconstruction, we initialize the world coordinate system with the first window and use the reconstructed frames (image and local point map produced by the I2P) as input for the L2W model. The L2W model incrementally register these local reconstructions into a unified global 3D coordinate system. To ensure both accuracy and efficiency during this process, the system maintains a limited reservoir of registered frames, called scene frames. Whenever the L2W model registers a new keyframe, we retrieve the best-correlated scene frames as a reference. The details are introduced in Sec. 3.2.
@@ 3.1 Inner-Window Local Reconstruction
The Image-to-Points (I2P) model aims to infer dense 3D pointmaps for every pixel of a keyframe in a given video clip. By default, the middle image of a window $W$ is chosen as the keyframe $I_{key}$ to define the local coordinate system, as it is most likely to have the largest overlap with other frames. The remaining images $\{I_{sup_{i}}\}_{i=1}^{L-1}$ serve as supporting frames. Note that the 3D pointmaps of supporting frames can also be reconstructed through I2P.
@@
Image encoder.
For a given video clip, the image encoder $E_{img}$ encodes each frame $I_{i}$ to obtain token representations $F_{i}\in\mathbb{R}^{T\times d}$, where $T$ is the number of tokens and $d$ is the token dimension. The encoder $E_{img}$ comprises $m$ ViT encoder blocks, each containing self-attention and feed-forward layers. The encoding process is denoted as
@@
Keyframe decoder.
The keyframe decoder $D_{key}$ consists of $n$ ViT decoder blocks, each containing self-attention, cross-attention, and feed-forward layers. Unlike DUSt3R which uses the standard cross-attention, we introduce a novel multi-view cross-attention to combine information from different supporting frames. Given the feature tokens $F_{key}$ and $\{F_{sup_{i}}\}_{i=1}^{L-1}$, the keyframe decoder $D_{key}$ takes $F_{key}$ as input for self-attention and performs cross-attention between $F_{key}$ and $\{F_{sup_{i}}\}_{i=1}^{L-1}$. A decoder block is illustrated in Figure 3. For each cross-attention layer, queries are taken from $F_{key}$, while keys and values are extracted from the supporting tokens $F_{sup_{i}}$. These $L-1$ cross-attention layers are independent of each other, allowing for parallel processing. A max-pooling layer is then employed to aggregate features after cross-attention. We obtain decoded keyframe tokens $G_{key}$ as:
@@
Supporting decoder.
The supporting decoder $D_{sup}$ is designed to complement the keyframe decoder. It inherits the decoder architecture used in DUSt3R, consisting of $n$ standard ViT decoder blocks. The cross-attention mechanism is applied only to exchange information with the keyframe. Note that all supporting frames share the same $D_{sup}$. This process is denoted as
@@
Points reconstruction.
Similar to DUSt3R, we apply a linear head [62] to regress dense 3D pointmaps in the unified coordinate system from decoded tokens. In addition to the pointmaps, we also predict the confidence maps for all frames to evaluate their reliability. The final predictions are:
@@
Training loss.
Following DUSt3R, the I2P network is trained end-to-end using ground-truth scene points $\{X_{i}\}_{i=1}^{L}$. Both the ground truth and predicted point maps are normalized to a canonical scale, determined by the average distance of all valid points within the window to the origin. The confidence-aware training loss is:
@@ 3.2 Inter-Window Global Registration
Scene initialization.
The first window is used to initialize the scene model. It’s crucial to ensure this initialization is as accurate as possible. To achieve this, we execute the I2P network $L$ times, attempting to traverse and designate each frame within the window as the keyframe. We then select the result with the highest total confidence score for scene model initialization. This process results in a scene pointcloud along with a set of registered frames. All these frames are regarded as scene frames, and they are used to initialize the buffering set.
Reservoir and retrieval.
Each scene frame $I_{sce_{i}}$ is recorded with its latent feature $F_{sce_{i}}$ and pointmaps $\tilde{X}_{sce_{i}}$. For efficiency, we apply reservoir sampling to allow storing an unbiased subset of an empirical distribution in a bounded amount of memory. The first $B$ registered frames chosen are directly inserted into the buffering set. For each subsequent frame with $id>B$, the probability of inserting it is $B/id$. If chosen for insertion, it will randomly replace one of the current scene frames in the buffering set.
Given a new keyframe $I_{key}$ to be registered, we feed its feature $F_{key}$ and the features from the buffering set into a retrieval module
$\text{Retrieval}(F_{key}^{(T\times d)},\{F_{sce_{i}}^{(T\times d)}\})$
@@
Points embedding.
The 3D pointmaps reconstructed by the I2P model are encoded into the L2W model using a patch embedding method similar to image patchification in the ViT encoder $E_{img}$. We process the new keyframe and $K$ retrieved scene frames in parallel as:
@@
This resulting a token set $\{\mathcal{F}_{key},\{\mathcal{F}_{sce_{i}}\}_{i=1}^{K}\}$ contains joint features of image patch appearance and 3D geometry for the keyframe and retrieved scene frames.
Registration decoder.
The registration decoder $D_{reg}$ takes feature tokens $\{\mathcal{F}_{key},\{\mathcal{F}_{sce_{i}}\}_{i=1}^{K}\}$ as input and aims to transform the local reconstruction of the keyframe to the scene coordinate system. It takes the same network architecture of the keyframe decoder $D_{key}$. This decoding process is denoted by
@@
Scene decoder.
The scene decoder $D_{sce}$ takes the token set $\{\mathcal{F}_{key},\{\mathcal{F}_{sce_{i}}\}_{i=1}^{K}\}$ as input to refine the scene geometry without coordinate system changes. It uses the same network architecture as the keyframe decoder $D_{key}$, allowing us to extend to multi-keyframe co-registration (see supplementary material for details). By default, we register one keyframe each time. Each of the $\mathcal{F}_{sce_{i}}$ has information exchange only with the $\mathcal{F}_{key}$. This decoding process is denoted by
@@
Points reconstruction and training loss.
We apply the same head design as that of the I2P network to predict all the pointmaps $\tilde{X}_{i}$ in the global scene coordinate system:

### v3

@@ Abstract
In this paper, we introduce SLAM3R, a novel and effective system for real-time, high-quality, dense 3D reconstruction using RGB videos. SLAM3R provides an end-to-end solution by seamlessly integrating local 3D reconstruction and global coordinate registration through feed-forward neural networks. Given an input video, the system first converts it into overlapping clips using a sliding window mechanism. Unlike traditional pose optimization-based methods, SLAM3R directly regresses 3D pointmaps from RGB images in each window and progressively aligns and deforms these local pointmaps to create a globally consistent scene reconstruction - all without explicitly solving any camera parameters. Experiments across datasets consistently show that SLAM3R achieves state-of-the-art reconstruction accuracy and completeness while maintaining real-time performance at 20+ FPS. Code available at: https://github.com/PKU-VCL-3DV/SLAM3R.
@@ 1 Introduction
Figure 1: We introduce a novel dense reconstruction system - SLAM3R. It takes a monocular RGB video as input and reconstructs the scene as a dense pointcloud. The video is converted into short clips for local reconstruction (denoted as inner-window), which are then incrementally registered together (inter-window) to create a global scene model. This process runs in real-time, producing a reconstruction that is both accurate and complete.
@@
While monocular dense SLAM systems encounter the limitations mentioned earlier, recent advances in two-view geometry have shown promising potential. DUSt3R [64] introduces a purely end-to-end approach for learning dense reconstruction. Trained on large-scale datasets, its network is capable of producing high-quality dense reconstructions from paired images in real-time. However, for multiple views, a global optimization step is required to align these image pairs, which significantly hampers its efficiency. A concurrent work, Spann3R [61], extends DUSt3R to multi-view (video) scenarios through an incremental pipeline with spatial memory. While this method accelerates the reconstruction process, it unfortunately results in noticeable accumulated drift and reduced reconstruction quality.
@@
  • We present a novel real-time end-to-end dense 3D reconstruction system that uses RGB videos to directly predict 3D pointmaps in a unified coordinate system through feed-forward neural networks.
@@ 2 Related Work
Traditional offline approaches. Dense 3D pointcloud reconstruction is a long-standing problem in computer vision. Classical approaches to this problem first determine camera parameters using Structure from Motion (SfM) [53, 66, 47, 31, 33], followed by dense 3D points triangulation with Multi-View Stereo (MVS) [18, 48, 68, 60, 1]. In recent years, neural implicit [36, 62, 30, 65, 8] and 3D Gaussian [19, 20, 12] representations have been applied to further enhance the quality of dense reconstruction. While these methods deliver high-quality results, they have a significant limitation: the requirement for offline processing to generate the final 3D model, which restricts their applicability in real-time scenarios. In this paper, we focus on online dense reconstruction in the context of Simultaneous Localization and Mapping (SLAM).
Dense SLAM. Early works on SLAM [15, 4, 25, 16, 38, 37, 7] focused on reconstructing the structure of unknown environments while simultaneously localizing camera poses. These approaches prioritize real-time performance but produce only sparse structures of the scene. Dense SLAM approaches [39, 76, 5, 10, 27, 56, 17, 55, 22, 78, 9, 42, 11] incorporate detailed scene geometry information to improve pose estimation. DROID-SLAM [56] introduces recurrent iterative updates of camera poses and pixel-wise depth estimates, while TANDEM [27] proposes an online MVS module for depth prediction. These systems enable real-time dense scene reconstruction. However, their focus on camera trajectory accuracy often results in incomplete and noisy 3D reconstruction. Neural implicit and Gaussian representations have also been integrated with dense SLAM systems [55, 22, 78, 42, 72, 32, 23, 67, 34, 21, 9, 42, 45, 44]. However, these approaches often rely on additional depth sensors or focus primarily on novel view synthesis rather than producing detailed geometric reconstruction.
@@
End-to-end dense 3D reconstruction. DUSt3R [64] introduces the first purely end-to-end dense 3D reconstruction pipeline without relying on camera parameters. Recently, several works have adopted a similar approach for single-view reconstruction [63], feature matching [28], novel view synthesis [52, 70], and dynamics reconstruction [74]. These successes demonstrate the effectiveness of end-to-end dense point prediction, inspiring us to develop a dense reconstruction system with a similar methodology.
While DUSt3R operates in real-time for two-view predictions, its extension to multiple views involves exhaustive pairing images and performing an additional global optimization step. This process significantly increases computational time, thereby hindering its real-time performance. MASt3R [28] enhances the matching capability of DUSt3R by adding a match head, achieving more accurate keypoint correspondences for 3D reconstruction [14], but at the cost of increased computational time. More recently, the concurrent work Spann3R [61] extends DUSt3R with spatial memory. It takes a video as input and performs incremental scene reconstruction in a unified coordinate system without requiring global optimization. While this approach significantly improves runtime efficiency, the incremental reconstruction pipeline frame by frame leads to noticeable accumulated drift. Unlike Spann3R, our networks at each hierarchy take multiple frames as input to minimize drift. Additionally, we propose a self-contained retrieval module that, when registering a new frame, this module selects not only its previous few frames but also other similar frames from long-term history for more global scene reference.
@@ 3 Method
Problem statement. Given a monocular video consisting of a sequence of RGB image frames $\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$ that captures a static scene, the goal is to reconstruct its dense 3D poincloud $P\in\mathbb{R}^{M\times 3}$, where $M$ is the number of 3D points. Our work focuses on three key objectives: maximizing 3D points recovery for reconstruction completeness, improving the accuracy of each recovered point, and achieving these goals while preserving real-time performance.
System overview. Figure 2 illustrates an overview of the proposed dense reconstruction system. It consists of two main components: an Image-to-Points (I2P) network that recovers local 3D points from video clips, and a Local-to-World (L2W) network that registers local reconstructions into a global scene coordinate system. During the reconstruction of the dense point cloud, the system does not explicitly solve any camera parameters. Instead, it directly predicts 3D point maps in unified coordinate systems.
The system starts by applying a sliding window mechanism of length $L$ to convert the input video into short clips $\{\mathcal{W}_{i}\in\mathbb{R}^{L\times H\times W\times 3}\}$. The I2P network then processes each window $\mathcal{W}_{i}$ to recover local 3D pointmaps. Within each window, the system selects a keyframe to define a reference coordinate system for point reconstruction, as detailed in Sec. 3.1. By default, the stride of the sliding window is set to 1, ensuring each input frame in the video is selected at least once as a keyframe. For global scene reconstruction, we initialize the world coordinate system with the first window and use the reconstructed frames (image and local point map produced by the I2P) as input for the L2W model. The L2W model incrementally register these local reconstructions into a unified global 3D coordinate system. To ensure both accuracy and efficiency during this process, the system maintains a limited reservoir of registered frames, called scene frames. Whenever the L2W model registers a new keyframe, we retrieve the best-correlated scene frames as a reference. The details are introduced in Sec. 3.2.
@@ 3.1 Inner-Window Local Reconstruction
The Image-to-Points (I2P) model aims to infer dense 3D pointmaps for every pixel of a keyframe in a given video clip. By default, the middle image of a window $\mathcal{W}$ is chosen as the keyframe $I_{key}$ to define the local coordinate system, as it is most likely to have the largest overlap with other frames. The remaining images $\{I_{sup_{i}}\}_{i=1}^{L-1}$ serve as supporting frames. Note that the 3D pointmaps of supporting frames can also be reconstructed through I2P.
@@
Image encoder. For a given video clip, the image encoder $E_{img}$ encodes each frame $I_{i}$ to obtain token representations $F_{i}\in\mathbb{R}^{T\times d}$, where $T$ is the number of tokens and $d$ is the token dimension. The encoder $E_{img}$ comprises $m$ ViT encoder blocks, each containing self-attention and feed-forward layers. The encoding process is denoted as
@@
Keyframe decoder. The keyframe decoder $D_{key}$ consists of $n$ ViT decoder blocks, each containing self-attention, cross-attention, and feed-forward layers. Unlike DUSt3R which uses the standard cross-attention, we introduce a novel multi-view cross-attention to combine information from different supporting frames. Given the feature tokens $F_{key}$ and $\{F_{sup_{i}}\}_{i=1}^{L-1}$, the keyframe decoder $D_{key}$ takes $F_{key}$ as input for self-attention and performs cross-attention between $F_{key}$ and $\{F_{sup_{i}}\}_{i=1}^{L-1}$. A decoder block is illustrated in Figure 3. For each cross-attention layer, queries are taken from $F_{key}$, while keys and values are extracted from the supporting tokens $F_{sup_{i}}$. These $L-1$ cross-attention layers are independent of each other, allowing for parallel processing. A max-pooling layer is then employed to aggregate features after cross-attention. We obtain decoded keyframe tokens $G_{key}$ as:
@@
Supporting decoder. The supporting decoder $D_{sup}$ is designed to complement the keyframe decoder. It inherits the decoder architecture used in DUSt3R, consisting of $n$ standard ViT decoder blocks. The cross-attention mechanism is applied only to exchange information with the keyframe. Note that all supporting frames share the same $D_{sup}$. This process is denoted as
@@
Points reconstruction. Similar to DUSt3R, we apply a linear head [64] to regress dense 3D pointmaps in the unified coordinate system from decoded tokens. In addition to the pointmaps, we also predict the confidence maps for all frames to evaluate their reliability. The final predictions are:
@@
Training loss. Following DUSt3R, the I2P network is trained end-to-end using ground-truth scene points $\{X_{i}\}_{i=1}^{L}$. Both the ground truth and predicted point maps are normalized to a canonical scale, determined by the average distance of all valid points within the window to the origin. The confidence-aware training loss is:
@@ 3.2 Inter-Window Global Registration
Scene initialization. The first window is used to initialize the scene model. It’s crucial to ensure this initialization is as accurate as possible. To achieve this, we execute the I2P network $L$ times, attempting to traverse and designate each frame within the window as the keyframe. We then select the result with the highest total confidence score for scene model initialization. This process results in a scene pointcloud along with a set of registered frames. All these frames are regarded as scene frames, and they are used to initialize the buffering set.
Reservoir and retrieval. Each scene frame $I_{sce_{i}}$ is recorded with its latent feature $F_{sce_{i}}$ and pointmaps $\tilde{X}_{sce_{i}}$. For efficiency, we apply reservoir sampling to allow storing an unbiased subset of an empirical distribution in a bounded amount of memory. The first $B$ registered frames chosen are directly inserted into the buffering set. For each subsequent frame with $id>B$, the probability of inserting it is $B/id$. If chosen for insertion, it will randomly replace one of the current scene frames in the buffering set.
Given a new keyframe $I_{key}$ to be registered, we feed its feature $F_{key}$ and the features from the buffering set into a retrieval module,
$\text{Retrieval}(F_{key}^{(T\times d)},\{F_{sce_{i}}^{(T\times d)}\}),$
@@
Points embedding. The 3D pointmaps reconstructed by the I2P model are encoded into the L2W model using a patch embedding method similar to image patchification in the ViT encoder $E_{img}$. We process the new keyframe and $K$ retrieved scene frames in parallel as:
@@
This resulting a token set $\{\mathcal{F}_{key},\{\mathcal{F}_{sce_{i}}\}_{i=1}^{K}\}$ contains joint features of image patch appearance and 3D geometry for the keyframe and retrieved scene frames. In the following decoders, $\{\mathcal{P}\}$ are further accumulated to ${}\{\mathcal{F}\}$ between adjacent blocks to enhance the geometric representation.
Registration decoder. The registration decoder $D_{reg}$ takes feature tokens $\{\mathcal{F}_{key},\{\mathcal{F}_{sce_{i}}\}_{i=1}^{K}\}$ as input and aims to transform the local reconstruction of the keyframe to the scene coordinate system. It takes the same network architecture of the keyframe decoder $D_{key}$. This decoding process is denoted by
@@
Scene decoder. The scene decoder $D_{sce}$ takes the token set $\{\mathcal{F}_{key},\{\mathcal{F}_{sce_{i}}\}_{i=1}^{K}\}$ as input to refine the scene geometry without coordinate system changes. It uses the same network architecture as the keyframe decoder $D_{key}$, allowing us to extend to multi-keyframe co-registration (see supplementary material for details). By default, we register one keyframe each time. Each of the $\mathcal{F}_{sce_{i}}$ has information exchange only with the $\mathcal{F}_{key}$. This decoding process is denoted by
@@
Points reconstruction and training loss. We apply the same head design as that of the I2P network to predict all the pointmaps $\tilde{X}_{i}$ in the global scene coordinate system: